OpenStudy (anonymous):

Alan can chop 3/8 of a carrot in 1/8 min. How many carrots can he chop in 1 min?

OpenStudy (unanimoose):

\[\frac{ 3 }{ 8 } \times \frac{ 8 }{ 1 } = x\]

OpenStudy (unanimoose):

Can you work that out first?

OpenStudy (unanimoose):

Ok well: \[\frac { 3 }{ 8 } \times \frac { 8 }{ 1 } = x = \frac {(3 \times 8)}{ 8 }\]

OpenStudy (unanimoose):

So multiply the numerators and divide them by 8, which will be the amount of carrots he can chop in 1 minute.